# Superconducting qubits at the utility scale: the potential and limitations of modularity

### v2

@@ II.2.2 Fermi-Hubbard Hamiltonian simulations
Of note is that in [73] it was determined that the time required to run these studies on a CPU varies like,
$t_{\text{C}PU}\propto L_{x}L_{y}(2\chi^{3}D^{4}d^{2}+\chi^{2}D^{2}d^{2}+\chi^{2}D^{4}d^{4}),$
where $\chi$ is the truncated bond dimension and $d$ the problem dimension. Memory scaled as,
$2\chi^{2}D^{4}d^{2}M_{\text{n}um}+2N\chi^{2}D^{2}dM_{\text{n}um}.$
The authors of the study empirically found $\chi=3D$ was sufficiently accurate for their studies, with error about $1\%$. These estimates provide an interesting guidepost for the tipping point at which FTQCs quantum computers may outperform classical computers for these types of problems.
II.2.3 Algorithm and Approach
Owing to its optimal scaling characteristics, in terms of queries to the block encoding of the Hamiltonian [14], we have chosen Quantum Signal Processing (QSP) as the Hamiltonian simulation algorithm for which we obtain a resource estimate. QSP is briefly described below, but more details can be found in [14], [78], [79], and [80].
Given a Hamiltonian $H$ and simulation time $t$, QSP performs a sequence of single qubit rotations, parameterized by a sequence of phase angles, interlaced with a quantum walk-like operator. This allows one to apply a polynomial approximation of $e^{-iHt}$ to the input target state. The sequence of phase angles is responsible for the polynomial approximation. It is computed classically offline, according to the error tolerance of the approximation error desired, captured by the user-defined parameter $\epsilon_{\text{qsp}}$. $\epsilon_{\text{qsp}}$ is also inversely related to the probability of success of the QSP algorithm. Thus, for an accurate simulation, with a high probability of success, $\epsilon_{\text{qsp}}$ may have to be set relatively small.
For implementation details, we follow the examples accompanying the pyLIQTR [81] package (v1.3.3), as well as the work done in [82, 83].
For demonstration purposes, we set the desired probability of success of our algorithm, $p_{\rm algo}$, to be equal to 0.95. In real world applications this is typically fixed to meet outcome requirements (such as accuracy, wall clock time, etc). Denoting the probability of success of a QSP circuit, $p_{\rm qsp}$, we therfore require,
$p_{\rm algo}\leq p_{\rm qsp}.$ (26)
Since circuit size is inversely related to $\epsilon_{\text{qsp}}$, it is prudent to choose $\epsilon_{\text{qsp}}$ to be as large as possible while still guaranteeing that $p_{\rm qsp}$ exceeds the algorithm target of 0.95. According to [14] $(1-2\epsilon_{\text{qsp}})\leq p_{\rm qsp}$. Thus, to ensure eq. 26 is satisfied we set,
$1-2\epsilon_{\text{qsp}}=p_{\rm algo},$ (27)
and so in this work we choose $\epsilon_{\text{qsp}}$ to be equal to 0.025.
II.3 Circuit Synthesis
For our Hamiltonian simulation circuits, we used pyLIQTR v1.3.3 to construct the Hamiltonians and synthesize the QSP circuits. The pyLIQTR package provides tools that construct the Hamiltonians, calculate the necessary phase angles for QSP simulation, and generate QSP circuits. In addition, pyLIQTR allows the user to request random phase angles rather than solving for the precise values required for an accurate simulation. In our work, we used random phase angles. This is appropriate because we only perform resource estimation, not numerically accurate simulations. The correct number of phase angles is used in the circuits, but their numerical values are random. This provides a conservative overestimate of the resource estimations without explicitly incurring the time-consuming step of solving for the phase angles.
For Fermi-Hubbard models, we utilized pyLIQTR’s FermiHubbard model class, and employed the FermiHubbardSquare block encoding, which is based on [84].
III Resource estimation results
Key physical resource requirements for the test cases described in section II, over various problem sizes, are summarized in table 2. We demonstrate that our tool can estimate resources for problem sizes consistent with studies in the literature and beyond the scale of current studies.
Our methodology allows us to extend the state of the art in a few ways. (1) A full accounting for resources for both T counts, but also for executing the Clifford portions of the circuit as well. (2) A breakdown of physical and temporal resources allocated to portions of the execution of the algorithm, namely: Clifford, T-state and teleportation between modules. (3) A quantitative estimation of the impact on total resources of the performance of subsystems.
When we visualize this ensemble of compiled logical resources and physical resource requirements (fig. 9 and fig. 10 respectively), we see the run time and ancilla bus size trajectories that show an increasing trend in resources with increasing problem size. This provides us with some validation of the reasonableness of our methodology and its benefit as a tool to support studies into the future. Interestingly, we show a decreasing trend in logical qubits for small-sized problems until the number of modules increases. Once the number of modules increases, we see a sharp increase in the number of logical qubits. This is because physical qubits are assigned to T factories instead of logical qubits or the ancilla bus. As the problem size grows, the number of qubits used in the bilinear quantum memory and ancilla bus pattern increases in the single-module regime. This reduces the number of T factories that can fit in the module. As the T factories are large, monolithic regions of logical qubits, removing them gives a net reduction in the number of logical qubits in use, even though the number of patches in the bilinear pattern is larger. The storage space occupied in a room can decrease even as smaller objects are added, provided those smaller objects require the removal of large objects to accommodate them.
Table 2: Rigetti’s resource estimation results on a superconducting fault-tolerant quantum computing architecture with space-optimized graph-state compilation. Here, the Logical Bus includes patches for the memory and the ancilla bus mediating logical measurements in the bilinear pattern of fig. 6. Likewise, $T$ Patches includes logical patches for all $T$ Distillation Factories and $T$ Buffers.
Input circuit Logical Resources Superconducting FTQC resources
Instance logical qubits Logical Bus T Patches Physical qubits Modules $t_{\text{tot}}^{\text{hardware}}$ $E_{\text{tot}}$
Fermi Hubbard
2x2 37 276 4.05k 1.7M 2 4.16s 679Wh
4x4 74 332 3.22k 1.82M 2 37.5s 6.12kWh
8x8 182 728 2.07k 1.81M 2 9.5m 93kWh
16x16 578 2.31k 1.25k 3.45M 4 6.9h 8.11MWh
20x20 878 3.52k 1.4k 5.21M 6 37.8h 66.7MWh
50x50 5.05k 20.2k 5.54k 37.6M 42 7.13y 772GWh
100x100 10.1k 40.7k 24k 141M 174 85.9ky 38.5TWh
@@ III Resource estimation results
2x2 67 636 2.43k 1.77M 2 64.4s 10.5kWh
5x5 212 2.42k 816 3.43M 4 6.75h 7.92MWh
10x10 654 10.1k 2.78k 17.4M 20 5.07d 715MWh
19x19 2.22k 41.1k 13.5k 92M 110 82.6d 64.1GWh
Figure 9: Fault-tolerant compiled logical qubits and T count for three sets of test cases. Grey indicates Quantum Fourier Transform results for sizes 4 to 1000. The black plot is the Hamiltonian evolution of the Transverse-Ising model on triangular lattice sizes ranging from 2x2 to 19x19. The teal line is estimates for Hamiltonian simulation of a Fermi-Hubbard model on a square lattice with sizes ranging from 2x2 to 100x100.
Figure 10: Fault-tolerant hardware runtime and required number of physical qubits in the bilinear pattern for three sets of test cases. Grey indicates Quantum Fourier Transform results for sizes 4 to 1000. The black plot is the Hamiltonian evolution of the Transverse-Ising model on triangular lattice sizes ranging from 2x2 to 19x19. The teal line is an estimate for the Hamiltonian simulation of a Fermi-Hubbard model on a square lattice with sizes ranging from 2x2 to 100x100. We set the Hamiltonian evolution time to 1 in units of $1/J$ for both the Transverse-Ising and Fermi-Hubbard models.
@@
Figure 11: Physical qubit allocations for the types of qubits described in the microarchitecture shown in fig. 1; those involved in T state distillation, logical register and ancillas for queuing and routing.
@@
Figure 12: Proportion of quantum computer compute time allocated to graph state consumption (including T state injection), intermodule communications, and decoding delay for time evolutions.
@@
Figure 13: Comparison of normalized fault-tolerant runtime and number of intermodule pipes. The number of intermodule pipes varies from 1 to 32. Run time is normalized to the value with one pipe. Black indicates Quantum Fourier Transform results of size 10. The grey plot is the Hamiltonian evolution of the Transverse-Ising model on a triangular lattice of size 10. The teal line captures estimates for Hamiltonian simulation of a Fermi-Hubbard model on a square lattice with size 10x10.
@@
Figure 14: Total physical qubits required for Hamiltonian evolution of the Fermi-Hubbard of various sizes. Line colors denote different decoders; Teal indicates Minimum Weight Perfect Matching (MWPM), black is belief propagation, grey is belief propagation with ordered statistics decoding, and magenta is Astra [85]
The full accounting of both Clifford and T contributions to total resources, (1), is shown in fig. 11 for all families of applications for which we estimated resources. It is interesting to note that conventional T-counting methodologies don’t often surface the competition for resources between the Clifford and T-state generation of a quantum architecture. As problems scale, physical qubits need to be allocated to the increased needs of Clifford processing, squeezing out available qubits that could be allocated to MSD. This, in turn, reduces the number of MSD factories that can be applied to a calculation and thus increases runtime in a non-linear fashion.
@@
This impact on the breakdown of total runtime, (2), can be seen in fig. 12. Here, the contribution to runtime of graph consumption (including T state injection and graph state consumption) increases dramatically for the largest applications considered. This is because fewer qubits are allocated to create higher precision T-states, doubling the penalty in runtime. The impact of runtime on the teleportation present in the distributed architecture can also be seen in this figure. Despite a relatively slow inter-module interaction time (5x slower than the native code cycle), the contribution to overall runtime is application dependent. For QFT and transverse Ising applications, a relatively small portion of the overall runtime is spent performing this intermodule interaction, even at large problem sizes. This shows that computation can be partitioned for specific architectures and algorithms to allow distributed computing, with relatively low impact. However, this seems algorithm-dependent, and for Fermi-Hubbard problems, the penalty paid for distributed quantum computation is high.
@@
Moreover, we can test the sensitivity of runtime to parameters in this architectural model, (3). In fig. 13 we select an instance from each family, vary the number of inter-module connections, and calculate the impact on runtime. We define this in terms of a pipe where the number of connections equals the required code distance for the algorithm. Additional pipes will allow the teleportation of nodes between widgets to happen in parallel. This is done to quantify the penalty being paid for not having a fully connected lattice of qubits between modules. Here we see that while there is indeed an increased runtime for a reduced number of inter-module connections, there are diminishing returns for adding additional connections. Once again, this impact is seen to be highly application-dependent.
@@
Similarly, we can directly quantify the impact of changing the decoders on the physical system size. Results for Fermi-Hubbard instances of increasing problem size are shown in fig. 14 and reflect the utility and flexibility of our resource estimation tool to asses the impacts of changing multiple subsystems within the overall architecture.
IV Future Work
@@
The results from Section III can be improved by focusing on at least the following two aspects: a) using a decoder which is faster and/or has better scaling than the standard choice of MPWM; b) optimizing the widget circuits by, for example, reducing circuit gate counts and depths. The savings potential of these two approaches will be detailed in the following.
IV.1 Replacing the MWPM decoder
@@
The MWPM decoder (e.g.,  [86]) is the de facto decoder for performing resource estimates. Other decoders have been proposed, but MWPM is still the most efficient regarding the decoding speed vs. physical error rate trade-off: MWPM can handle high error rates while its complexity is polynomial. There are faster, lower polynomial complexity decoders (e.g., union-find variations or pure belief-propagation) and decoders with linear complexity, which are challenging to scale to high distances (machine learning decoders). Nevertheless, to the best of our knowledge, no alternative decoder exists that has, at the same time, a lower degree of polynomial decoding complexity and a threshold higher than MWPM.
@@
Herein, we report results, fig. 15, about a machine learning decoder that outperforms MWPM: it has a higher threshold and linear runtime complexity. Our machine learning decoder, Astra [85], uses graph neural networks (GNNs) in a novel way: we are learning a message-passing algorithm on the Tanner graph of the codes. We can train, within hours, such decoders with gaming GPUs for surface codes up to distance 11. In contrast to other machine learning decoders, the training time of our decoder is significantly lower (e.g. [87]) for comparable depths.
@@
Figure 15: The Logical Error Rate (LER) for code capacity depolarising noise of Astra vs MWPM. Astra has a threshold of $\sim 17\%$, and MWPM has a threshold of $\sim 14\%$. Astra clearly outperforms MWPM in terms of LER. In fact, Astra d9’s results are better than MWPM d11’s.
@@
We train distinct GNN decoders for each distance using syndromes collected by passing the surface codes through a code capacity noise channel. This channel assumes that the data qubits are affected by noise with a probability $p$ and that the syndrome measurements are perfect. For this work, we do not report the decoding of circuit-level correlated noise (e.g., [86]) and focus entirely on correcting depolarizing noise.
@@
Usually, the performance scaling equations for the surface code are obtained by simulations with circuit level noise (e.g., eq. 1) because this is considered the most exact model. Nevertheless, the code capacity is more straightforward for benchmarking the preliminary decoder versions. To allow the comparison between MWPM and Astra, we also benchmark MWPM with code capacity noise. The results are illustrated in Fig. 15. Based on the MWPM code capacity errors, we obtain the MWPM scaling equation as
@@
We observe that Astra has a higher threshold compared to MWPM ($17\%>14\%$) and offers faster (better) logical error suppression ($0.56>0.52$). This implies that, by replacing the MWPM with Astra, one can achieve the same logical error rate with lower distance surface codes. After benchmarking its performance under circuit-level noise, future work will focus on quantifying the exact savings when utilizing the new decoder.
IV.2 Ultra-large scale circuit optimization
@@
Widgets (Section A.2.4) are subcircuits representing relevant parts of the larger circuits. Due to their size, widgets can be more easily analyzed and optimized. Nevertheless, even widgets are complicated to optimize for metrics such as $T$-count and $T$-depth. Optimization is a highly complex task, and most optimization heuristics are not designed (or do not perform as expected) for (sub)circuits of tens of qubits and hundreds of gates.
@@
Pandora offers a simple API that can be used to develop new and more complex optimization methods. We interfaced our tool with Google Qualtran [89] and can easily extract large circuits for analysis and optimization. Currently, the circuits and widgets available in Qualtran have been optimized manually, so automatic optimization is even more challenging. Therefore, the results from Table 3 belong to far from optimal circuits.
Table 3 reports preliminary results on optimizing arithmetic widgets as presented in [90]. The latter are ubiquitous building blocks of practical application circuits like the Fermi-Hubbard. Our optimization procedure for optimizing the circuits is not as complex as the one described by Ref. [90]. We implemented a heuristic that uses multiple threads for applying circuit template rewrites [90]: 1) decomposition of gates (e.g., Toffoli); 2) canceling H, T, X gates; 3) canceling CNOT gates; 4) commuting gates to the left and the right; 5) reversing directions of CNOT gates.
@@
V Conclusion
This paper proposes a fault-tolerant architecture based on modular quantum processing units utilizing superconducting qubits and the graph state formalism. We provide quantitative estimates for the scale and runtime of executing various quantum algorithms at the utility scale, employing a resource estimation framework and software tool. This framework estimates the physical resources required for quantum algorithms and offers estimates for test cases derived from existing literature. We have demonstrated that a fault-tolerant quantum computer built on a distributed superconducting architecture, consisting of two legs of modules, could support scientifically interesting applications within a reasonable runtime. However, a significant algorithm-dependent penalty exists for distributing computation across multiple modules. These estimates align with established results, such as factoring [2, 6]. This model-driven approach delineates how a superconducting platform could accomplish this, including details on scheduled execution over the preparation and consumption of graph states and hardware-level justifications for some of the most critical size and power considerations.
This work has also enabled research toward ensuring that real quantum computing platforms can be developed to provide the advantages of fault-tolerant operation. It serves as a software testbed for evaluating the impact of algorithmic, compilation, decoding, and physical-layer intrinsic or system architecture proposals. Moreover, this allows for direct per-algorithm comparison of different modalities by assessing run times, spatial, and energy resource costs. Rapid iteration in a simulated environment is a proven method for accelerating technology growth, and we hope that our efforts to produce this tooling contribute positively to this endeavor.
VI Acknowledgements

### v3

@@ IV Resource estimation results
[…]
@@
2x2 67 652 2.43k 1.41M 2 64.4s 840W 168nW 10.5kWh
5x5 212 2.98k 816 4.39M 6 6.64h 2.52kW 5.04nW 6.42MWh
[…]
Key physical resource requirements for the test cases described in section˜III across various problem sizes are summarized in table˜2. We demonstrate that our tool, RRE, can estimate resource requirements for problem sizes consistent with the literature and that extend beyond the scales of current studies.
Our methodology enables us to extend the state of the art in several ways. (1) A full accounting for resources for both T counts, but also for executing the Clifford portions of the circuit as well. (2) A breakdown of physical and temporal resources allocated to portions of the execution of the algorithm, namely: Clifford, $T$-state, and teleportation between modules. (3) A quantitative estimation of the impact on total resources of the performance of subsystems.
When we visualize this ensemble of compiled logical resources and physical resource requirements (fig.˜7 and fig.˜8, respectively), we observe trajectories of runtime and auxiliary bus size that increase with problem size. This provides validation of the reasonableness of our methodology and its utility as a tool to support future studies. Interestingly, we observe a decreasing trend in the number of logical qubits for small-sized problems until the number of modules increases. As the number of modules increases, we observe a sharp rise in the number of logical qubits. This is because physical qubits are assigned to $T$-factories rather than to logical memory qubits or the auxiliary bus. As the problem size grows, the number of physical qubits used in the bilinear quantum memory and auxiliary bus pattern increases in the single-module regime. This reduces the number of $T$-factories that can be accommodated within the module. As the $T$-factories are large monolithic regions of logical qubits, removing them gives a net reduction in the number of logical qubits in use, even though the number of patches in the bilinear pattern is larger. The available storage space in a room can decrease even as smaller objects are added, provided that those smaller objects require removing larger objects to accommodate them.
The full accounting of both Clifford and $T$-state contributions to total resources, (1), is shown in fig.˜9 for all families of applications for which we estimated resources. It is interesting to note that conventional $T$-counting methodologies do not often reveal resource competition between the Clifford and $T$-state generations of a quantum architecture. As problems scale, physical qubits must be allocated to meet the increased requirements of Clifford processing, thereby reducing the number of available physical qubits for MSD. This, in turn, reduces the number of $T$-factories that can be applied to a calculation, thereby increasing runtime in a non-linear fashion.
Figure 7: Total allocated logical qubits (in the bilinear pattern) and T-count for three sets of test cases. Grey indicates Quantum Fourier Transform results for sizes 4 to 1000. The black curve shows the Hamiltonian evolution of the Transverse-Ising model on triangular lattices ranging from 2x2 to 19x19. The teal line shows estimates from Hamiltonian simulations of a Fermi-Hubbard model on a square lattice, with sizes ranging from 2x2 to 100x100.
Figure 8: Total allocated physical qubits (in the bilinear pattern) and FT hardware runtime for three sets of test cases. Grey indicates Quantum Fourier Transform results for sizes 4 to 1000. The black curve shows the Hamiltonian evolution of the Transverse-Ising model on triangular lattices ranging from 2x2 to 19x19. The teal line is an estimate for the Hamiltonian simulation of a Fermi-Hubbard model on a square lattice with sizes ranging from 2x2 to 100x100. We set the Hamiltonian evolution time to 1 in units of $1/J$ for both the Transverse-Ising and Fermi-Hubbard models.
@@
Figure 9: Physical qubit allocations for the types of logical qubits described in the logical micro-architecture shown in fig.˜1; those involved in the logical memory register, auxiliary buses (for single-qubit measurements, queuing, and routing), the $T$-transfer bus, and $T$-factories.
@@
Figure 10: Proportion of FTQC’s compute time allocated to graph state consumption (using distilled $T$-states), intermodule communications, distillation, and decoding delays for time evolution problems.
@@
Figure 11: Comparison of normalized total FT runtime and number of intermodule pipes. The number of intermodule pipes varies from 1 to 32. Runtime is normalized to the corresponding value for one pipe. Black indicates Quantum Fourier Transform results of size 10. The grey plot is the Hamiltonian evolution of the Transverse-Ising model on a triangular lattice of size 10. The teal line represents estimates for the Hamiltonian simulation of a Fermi-Hubbard model on a 10x10 square lattice.
@@
Figure 12: Total number of allocated physical qubits (parameter 35 of appendix˜A) for Hamiltonian evolution of the Fermi-Hubbard of various sizes. Line colors denote different decoders; Teal indicates MWPM, black is the Belief Matching (BM) decoder, grey is Belief Propagation with ordered statistics decoding (BP-OSD), and magenta is Astra [45].
@@
This impact on the breakdown of total runtime, (2), is evident in fig.˜10. Here, the contribution to graph consumption runtime (using distilled $T$-states) increases dramatically for the largest applications considered. This is because fewer physical qubits are allocated to create higher precision $T$-states, increasing the runtime penalty. The impact of runtime on teleportation in the distributed architecture is also evident in this figure. Despite a relatively slow inter-module interaction time (5x slower than the native code cycle), the contribution to overall runtime is application dependent. For QFT and transverse Ising applications, a relatively small portion of the overall runtime is spent performing this intermodule interaction, even at large problem sizes. This shows that computation can be partitioned across specific architectures and algorithms to enable distributed computing with relatively low impact. However, this appears to be algorithm-dependent, and for Fermi-Hubbard problems, the penalty for distributed quantum computation is high.
@@
Moreover, we can test the sensitivity of runtime to parameters in this architectural model, (3). In fig.˜11, we select an instance from each family, vary the number of inter-module connections, and calculate the impact on runtime. We define this in terms of a pipe where the number of connections equals the required code distance for the algorithm. Additional pipes will enable parallel node-to-node teleportation between widgets. This is done to quantify the penalty associated with not having a fully connected lattice of physical qubits between modules. Here, we observe that while runtime increases with fewer inter-module connections, the marginal benefit of adding more connections diminishes. Once again, this impact is highly application-dependent.
@@
Similarly, we can directly quantify the impact of changing the decoders on the physical system size. Results for Fermi-Hubbard instances of increasing problem size are shown in fig.˜12 and demonstrate the utility and flexibility of our resource estimation tool for assessing the effects of changes across multiple subsystems within the overall architecture.
V Future work
@@
The results from Section IV can be improved by focusing on at least the following two aspects: a) using a decoder that is faster and may have better scaling than the standard choice of MPWM; b) optimizing the widget circuits by, for example, reducing circuit gate counts and depths. Below, we present the savings potential of these two approaches, along with additional details on the design of high-cooling-power cryostats.
V.1 Replacing the MWPM decoder
@@
The MWPM decoder (e.g., [62]) is the de facto standard for resource estimation. Other decoders have been proposed, but MWPM remains the most efficient with respect to the decoding speed-physical error rate trade-off: it can handle high error rates while maintaining polynomial complexity. There are faster, lower-polynomial-complexity decoders (e.g., union-find variants or pure belief propagation) and decoders with linear complexity, which are challenging to scale to high distances (e.g., machine learning decoders). Nevertheless, to the best of our knowledge, no alternative decoder exists that has, at the same time, a lower degree of polynomial decoding complexity and a threshold higher than MWPM.
@@
Herein, we report results in fig.˜13 for a machine-learning decoder that outperforms MWPM: it has a higher threshold and linear runtime complexity. Our machine learning decoder, Astra [45], uses graph neural networks (GNNs) in a novel way: we learn a message-passing algorithm on the Tanner graph of the code. We can train such decoders within hours on gaming GPUs for surface codes at distances up to 11. In contrast to other machine learning decoders, the training time of our decoder is significantly lower (e.g. [85]) for comparable depths.
@@
Figure 13: The Logical Error Rate (LER) for code capacity depolarising noise of Astra vs MWPM. Astra has a threshold of $\sim 17\%$, and MWPM has a threshold of $\sim 14\%$. Astra clearly outperforms MWPM in terms of LER. In particular, Astra d9’s results are better than those of MWPM d11.
@@
We train distinct GNN decoders for each distance using syndromes generated by passing the surface codes through a code-capacity noise channel. This channel assumes that the physical data qubits are affected by noise with a probability $p$ and that the syndrome measurements are perfect. For this work, we do not report the decoding of circuit-level correlated noise (e.g., [62]) and focus entirely on correcting depolarizing noise.
@@
Typically, the performance-scaling equations for the surface code are obtained from simulations with circuit-level noise (e.g., eq.˜1) because this is considered the most accurate model. Nevertheless, the code capacity is more straightforward to benchmark for preliminary decoder versions. To enable comparison with Astra, we also benchmark MWPM with code capacity noise. The results are illustrated in Fig. 13. Based on the MWPM code capacity errors, we obtain the MWPM scaling equation as
@@
We observe that Astra has a higher threshold than MWPM ($17\%>14\%$) and provides faster (better) logical error suppression ($0.56>0.52$). This implies that by replacing the MWPM with Astra, one can achieve the same logical error rate with lower-distance surface codes. After benchmarking its performance under circuit-level noise, future work will focus on quantifying the exact savings achieved by using the new decoder.
V.2 Ultra-large scale circuit optimization
@@
Widgets (detailed in section˜B.0.4) are subcircuits representing relevant parts of the larger circuits. Due to their size, widgets can be more easily analyzed and optimized. Nevertheless, even widgets are complicated to optimize for metrics such as T-count and $T$-depth. Optimization is a highly complex task, and most optimization heuristics are not designed (or do not perform as expected) for (sub)circuits of tens of logical qubits and hundreds of gates.
@@
Pandora provides a simple API for developing new and more complex optimization methods. We interfaced our tool with Google Qualtran [27] and can easily extract large circuits for analysis and optimization. Currently, the circuits and widgets available in Qualtran have been manually optimized; therefore, automatic optimization is even more challenging. Therefore, the results from Table 3 are far from optimal circuits.
Table 3 reports preliminary results on optimizing arithmetic widgets as presented in [55]. The latter are ubiquitous building blocks of practical application circuits, such as the Fermi-Hubbard model. Our circuit optimization procedure is not as complex as the one described in ref. [55]. We implemented a heuristic that uses multiple threads for applying circuit template rewrites [55]: 1) decomposition of gates (e.g., $TOFFOLI$); 2) canceling $H$, $T$, and $X$ gates; 3) canceling $CNOT$ gates; 4) commuting gates to the left and the right; 5) reversing directions of $CNOT$ gates.
@@
V.3 Cryostat design
The detailed design of high-cooling-power cryostats for operation at this scale is a significant challenge that will require dedicated effort over the coming years. For the resource estimates in this paper, we have produced total dissipation power (TDP) estimates for both the 4 K and 20 mK stages. These are presented in columns $P_{\text{Diss},\text{4K}}$ and $P_{\text{Diss},\text{20mK}}$ in table˜2, respectively. While the numbers appear daunting by today’s standards, several facilities provide well in excess of kW of cooling power at 2 K, such as the Linear Coherent Light Source II (LCLS II) at the Stanford Linear Accelerator Center (SLAC). This system can produce 4 kW of cooling power at 2 K and 18.5 kW at 4.5 K. While we consider a design for such a system beyond the scope of the current work, we believe it is feasible within the timescales required to develop it.
VI Conclusion
This paper proposes a fault-tolerant architecture based on modular quantum processing units that uses superconducting qubits, the surface code, and coherent inter-module connections. We provide quantitative estimates for the scale and runtime of various quantum computations. Our software tool compiles large-scale quantum circuits into a graph state, schedules them on a model of our logical architecture, and, from this, estimates the physical resources and time required for test cases derived from the existing literature. This model-driven approach delineates how a superconducting platform could accomplish the target computation, including details on the scheduled execution of graph state preparation and consumption, as well as hardware-level justifications for some of the most critical size and power considerations. We have demonstrated that such a fault-tolerant quantum computer could perform scientifically interesting computations, such as a $20\times 20$ Fermi-Hubbard time evolution, using 5.2M physical qubits in less than two days. However, the value of such a computation remains an open question. When utilized in the context of a complete high-temperature superconductivity study, the number of computations and their probability of success increase the estimated computation time to the order of years [3]. More work is needed to co-optimize this FTQC architecture with algorithms of interest, with distributed computation across multiple modules presenting as the major bottleneck in this study.
This work has enabled research toward the development of real quantum computing platforms that provide the advantages of fault-tolerant operations. It serves as a software testbed for evaluating the impact of intrinsic algorithmic, compilation, decoding, and physical-layer proposals, as well as system-architecture proposals. Moreover, this allows for direct per-algorithm comparison of different modalities by assessing run times, spatial, and energy resource costs. Rapid iteration in a simulated environment is a proven method for accelerating technological development, and we hope that our efforts to develop this tooling contribute positively to this endeavor.
VII Acknowledgments